### v2

# On biquandle-based invariant of immersed surface-links and on Yoshikawa oriented fifth move

Abstract.
We resolve an open problem showing that the Yoshikawa’s fifth oriented move in his list cannot be reproduced by any finite sequence of the other nine moves and planar isotopies. Our proof introduces a link-type semi-invariant that remains unchanged under all moves except the fifth; contrasting values on two equivalent diagrams force the move’s independence. Second, we extend the algebraic toolkit for immersed surface-links. After revisiting the banded-unlink description of immersed surfaces and the twelve local moves that relate their diagrams, we develop a coloring theory based on biquandles. By assigning elements of a biquandle to diagram arcs according to local rules, we obtain a counting invariant of immersed surfaces up to isotopy.
Key words and phrases:
triple-crossing diagram
2020 Mathematics Subject Classification:
57K45 (primary), secondary: 57Q35, 57R42, 57K12
1. Introduction
The paper tackles two central problems in the diagrammatic study of oriented surfaces immersed in $\mathbb{R}^{4}$.
First, it settles an open question about Yoshikawa’s moves, the ten local moves that generate all diagrammatic isotopies of oriented surface-links. Previous work had proved the independence of every move except the oriented fifth move. We construct a semi-invariant derived from admissible marked-graph diagrams whose value is unchanged by every move in the generating set of other moves. Exhibiting two diagrams of the same surface-link with different values.
Second, the paper enriches the algebraic toolkit for immersed (i.e. generically self-intersecting) surface-links. We revisit the singular marked-graph (or “banded unlink”) presentation, extending the catalog of oriented planar moves so that any two diagrams of equivalent immersed surfaces differ by the twelve types of moves. We equip an immersed surface-link with a biquandle structure. Each semi-arc of a singular marked diagram is colored by elements of a biquandle according to local rules that respect crossings, markers and singular points. Because the biquandle axioms are tuned to the $\Gamma$-moves, the number of colorings is an invariant of the underlying surface. The authors compute this new biquandle coloring invariant for an illustrative example, obtaining two admissible colorings with a four-element biquandle.
Taken together, the independence proof sharpens our understanding of the move calculus, while the coloring invariant supplies fresh algebraic data for distinguishing immersed surfaces, advancing both the combinatorial and algebraic sides of four-dimensional knot theory.
This paper is organized as follows. In Section 2, we will prove an open problem of the independence of Yoshikawa oriented fifth move $\Gamma_{5}$ from the other generating set of oriented moves $\Gamma_{1},\ldots,\Gamma_{8},\Gamma_{4}^{\prime},\Gamma_{6}^{\prime}$ connecting diagrams of isotopic oriented surface-links.
In Section 3 we review the singular marked graph diagram method of presenting immersed surface-links and present the oriented planar moves between these diagrams.
In Section 4 we present the biquandle structure of an immersed surface-link with a method to obtain a biquandle coloring from a singular marked graph diagram and present a new invariant for oriented immersed surface-links. We show an example of the calculation.
Acknowledgements
”This research was funded in whole or in part by NCN 2023/07/X/ST1/00157. For the purpose of Open Access, the author has applied a CC-BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission.”
2. Independence of the Yoshikawa oriented fifth move
An embedding (or its image when no confusion arises) of a closed (i.e. compact, without boundary) surface $F$ into the Euclidean $\mathbb{R}^{4}$ (or into the $\mathbb{S}^{4}=\mathbb{R}^{4}\cup\{\infty\}$) is called a surface-link (or surface-knot if it is connected).
Two surface-links are equivalent if there exists an orientation preserving homeomorphism of the four-space $\mathbb{R}^{4}$ to itself (or equivalently auto-homeomorphism of the four-sphere $\mathbb{S}^{4}$), mapping one of those surfaces onto the other. See [Kam17] for an introductory material on surface-links.
A marked graph diagram is a planar $4$-valent graph embedding, with the vertices decorated either by a classical crossing or marker decoration.
Any abstractly created marked graph diagram is an admissible diagram if and only if both its resolutions are trivial classical link diagrams.
It is known that the set of ten types of moves $\mathcal{G}=\{\Gamma_{1},\ldots,\Gamma_{8},\Gamma_{4}^{\prime},\Gamma_{6}^{\prime}\}$ (presented in Fig. 1 and  2), called oriented Yoshikawa moves ([Yos94]), is a generating set of moves that relates two marked graph diagrams (modulo a planar isotopy) presenting equivalent oriented surface-links. In ([JKL13], [JKL15]) it is shown that any Yoshikawa move from the set $\mathcal{G}\backslash\{\Gamma_{5},\Gamma_{8}\}$ is independent of the other nine types. It is an open problem (see [JKL15], [Oht16]) whether the move $\Gamma_{5}$ is independent of the other Yoshikawa moves from the set $\mathcal{G}$.
{lpic}
[…]
Figure 1. Moves, part I.
{lpic}
[…]
Figure 2. Moves, part II.
The independence of every Yoshikawa move will lead us to obtain a minimal generating set of moves on oriented graph diagrams. We will prove the following main theorem.
Theorem 2.1.
The Yoshikawa move $\Gamma_{5}$ cannot be realized by a finite sequence of Yoshikawa moves of the other nine types from the set $\mathcal{G}$, and planar isotopy.
Proof of theorem 2.1.
We define a semi-invariant $L_{*}$ such that it preserves its values after performing each move from the set $\mathcal{G}\setminus\{\Gamma_{5}\}$, and we construct two pairs of admissible diagrams $D_{1},D_{2}$ of equivalent surface-links such that $L_{*}(D_{1})\not=L_{*}(D_{2})$.
Define $L_{*}$ as an ambient isotopy class of classical links obtained from admissible diagrams by changing all oriented marker decorations as shown in Figure 3 respectively. It is straightforward to see that $L_{*}$ is unchanged by any move in $G\setminus\{\Gamma_{5}\}$.
{lpic}
[…]
Figure 3. Defining a semi-invariant $L_{*}$.
The diagrams $D_{1}$ and $D_{2}$ represent equivalent surface-links, with the desired property explained above, are illustrated in Figure 4, where we have that $L_{*}(D_{1})$ is ambient isotopy types of the $6_{1}$ knot (i.e. a nontrivial classical knot with the Alexander polynomial $2-5t+2t^{2}$), on the other hand, $L_{*}(D_{2})$ is ambient isotopy type of the trivial classical link.
{lpic}
[…]
Figure 4. The diagrams $D_{1}$ and $D_{2}$.
3. Singular banded unlinks
Let $X,Y$ be smooth ($C^{\infty}$) manifolds. Let $f:X^{n}\to Y^{m}$ be a smooth map. It is called an immersion if at each point $x\in X$ the induced differential is a monomorphism. By the Whitney immersion theorem, any smooth map $f:X^{n}\to Y^{m}$ can be approximated homotopically with arbitrary accuracy by an immersion when $m\geq 2n$.
We consider smooth immersions $f:X^{n}\to Y^{2n}$ such that the following three conditions are satisfied:
(i) #$\#|f^{-1}(f(x))|\leq 2$, (ii) there are only a finite number of points with #$\#|f^{-1}(f(x))|=2$, (iii) at each singularity $p=f(x)=f(y)$, there is a coordinate chart around $p$ where the two coordinate subspaces $\mathbb{R}^{n}\times 0$ and $0\times\mathbb{R}^{n}$ are exactly the immersed images of $f$ near $x$ and $y$ respectively. That is the map is ”self-transverse”.
By general position theorems for maps, any smooth map $f:X^{n}\to Y^{2n}$ can be approximated homotopically with arbitrary accuracy by an immersion described above.
When we consider the case of immersions $\mathbb{S}^{1}\to\mathbb{R}^{2}$ we have the ”classical” case, which includes, for example, planar projections of knots and links embedded in $\mathbb{R}^{3}$, loosing the (codimension two) knotting phenomena. The next-dimension case is $f:X^{2}\to\mathbb{R}^{4}$ when we have both the finite point generic intersections and nontrivial generic embeddings (a known generalization of the classical knot theory).
An immersion in this dimensions (or its image when no confusion arises) of a closed (i.e. compact, without boundary) surface $F$ into the Euclidean $\mathbb{R}^{4}$ (or into the $\mathbb{S}^{4}=\mathbb{R}^{4}\cup\{\infty\}$) is called a immersed surface-link (or immersed surface-knot if it is connected).
Two immersed surface-links are equivalent if there exists an orientation preserving homeomorphism of the four-space $\mathbb{R}^{4}$ to itself (or equivalently auto-homeomorphism of the four-sphere $\mathbb{S}^{4}$), mapping one of those surfaces onto the other. Fix an immersed surface-link $F$ embedded in a manifold $\mathbb{S}^{4}$. For an open neighborhood, denoted $N(F)$, the exterior of $F$ is $E(F):=\mathbb{S}^{4}\backslash N(F)$.
If two singular surface-links are equivalent, then their exteriors are diffeomorphic.
A singular link $L$ in $\mathbb{R}^{3}$ is the image of an immersion in the classical case $\iota:S^{1}\sqcup\cdots\sqcup S^{1}\rightarrow\mathbb{R}^{3}$ which is injective except at isolated double points that are not tangencies. At every double point $p$ we include a small disk $v\cong D^{2}$ embedded in $\mathbb{R}^{3}$. We refer to these disks as the vertices of $L$. The double points of a singular link $L$ correspond to the isolated double points of an immersed surface in $\mathbb{R}^{4}$.
For each vertex $v$ of $L$, these two opposite push-offs form a bigon in a neighborhood of $v$, which bounds an embedded disk $D_{v}$. This disk $D_{v}$ can be chosen so that its interior intersects $L$ transversely in a single point near $v$. For each vertex $v$ select such a disk $D_{v}$ (ensuring that all of these disks are pairwise disjoint).
{lpic}
[…]
Figure 5. A vertex of a marked singular link and the corresponding surface cuts.
Let $D_{L}$ denote the union of all of these embedded disks. Let $(L,\sigma,B)$ be an singular link with bands $B=\{b_{1},\dots,b_{n}\}$ and $\Delta_{1},\ldots,\Delta_{a}\subset\mathbb{R}^{3}$ be mutually disjoint $2$-disks with $\partial(\cup_{j=1}^{a}\Delta_{j})=L_{B+}$, and let $\Delta_{1}^{\prime},\ldots,\Delta_{b}^{\prime}\subset\mathbb{R}^{3}$ be mutually disjoint $2$-disks with $\partial(\cup_{k=1}^{b}\Delta_{k}^{\prime})=L_{-}$.
We define $\Sigma\subset\mathbb{R}^{3}\times\mathbb{R}=\mathbb{R}^{4}$ an immersed surface-link corresponding to $(L,\sigma,B)$ by the following cross-sections.
$(\mathbb{R}^{3}_{t},\Sigma\cap\mathbb{R}^{3}_{t})=\left\{\begin{array}[]{ll}(\mathbb{R}^{3},\emptyset)&\hbox{for $t>1$,}\\ (\mathbb{R}^{3},L_{B+}\cup(\cup_{j=1}^{a}\Delta_{j}))&\hbox{for $t=1$,}\\ (\mathbb{R}^{3},L_{B+})&\hbox{for $0<t<1$,}\\ (\mathbb{R}^{3},L_{+}\cup(\cup_{i=1}^{n}b_{i}))&\hbox{for $t=0$,}\\ (\mathbb{R}^{3},L_{+})&\hbox{for $-1/2<t<0$,}\\ (\mathbb{R}^{3},L_{-}\cup D_{L})&\hbox{for $t=-1/2$,}\\ (\mathbb{R}^{3},L_{-})&\hbox{for $-1<t<-1/2$,}\\ (\mathbb{R}^{3},L_{-}\cup(\cup_{k=1}^{b}\Delta_{k}^{\prime}))&\hbox{for $t=-1$,}\\ (\mathbb{R}^{3},\emptyset)&\hbox{for $t<-1$.}\\ \end{array}\right.$
By an ambient isotopy of $\mathbb{R}^{3}$, we shorten the bands of a singular link with bands $LB$ so that each band is contained in a small $2$-disk. Replacing the neighborhood of each band with the neighborhood of a $4$-valent marked vertex as in Fig. 6, we obtain a singular marked graph.
{lpic}
[…]
Figure 6. A band corresponding to a marked vertex.
A singular marked graph diagram is a planar $4$-valent graph embedding, with the vertices decorated either by a classical crossing, marker or a singular decoration.
{lpic}
[…]
Figure 7. A neighborhood in the four-space of a marker and a singular point.
{lpic}
[]michal_14(3cm)
Figure 8. Example of an admissible singular marked graph diagram.
Any abstractly created singular marked graph diagram is an admissible diagram if and only if both its resolutions $L_{-}$ and $L_{B+}$ are trivial classical link diagrams.
Theorem 3.1 ([HKM21, Jab22c]).
Two oriented immersed surface-links $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ are equivalent if and only if their singular marked diagrams are related by a finite sequence of moves $\Gamma_{1},\ldots,\Gamma_{12}d$ shown in Figures 12, 911.
{lpic}
[…]
Figure 9. Moves, part III.
{lpic}
[…]
Figure 10. Moves, part IV.
{lpic}
[…]
Figure 11. Moves, part V.
4. Biquandle structure
A generalization of quandles (called biquandles) was introduced in [KauRad03]. A biquandle is an algebraic structure with two binary operations satisfying certain conditions which can be presented by semi-arcs of links (or semi-sheets of surface-links) as its generators modulo oriented Reidemeister moves (or Roseman moves). In [CES04], J. S. Carter, M. Elhamdadi, and M. Saito introduced and used cocycles to define invariants via colorings of link diagrams by biquandles and a state-sum formulation.
In [KKKL18] S. Kamada, A. Kawauchi, J. Kim and S. Y. Lee discussed the (co)homology theory of biquandles and developed the biquandle cocycle invariants for oriented surface-links by using broken surface diagrams generalizing quandle cocycle invariants. Then showed how to compute the biquandle cocycle invariants from marked graph diagrams.
Definition 4.1.
Let $X$ be a set. A biquandle structure on $X$ is a pair of maps ¯,¯$\,\underline{\triangleright}\,,\,\overline{\triangleright}\,:X\times X\to X$ satisfying:
  • (i)
    for all $x\in X$, $x\,\underline{\triangleright}\,x=x\,\overline{\triangleright}\,x$,
  • (ii)
    the maps $\alpha_{y},\beta_{y}:X\to X$ for all $y\in X$ and $S:X\times X\to X\times X$ defined by
    $\alpha_{y}(x)=x\,\overline{\triangleright}\,y,\quad\beta_{y}(x)=x\,\underline{\triangleright}\,y\quad\mathrm{and}\quad S(x,y)=(y\,\overline{\triangleright}\,x,x\,\underline{\triangleright}\,y)$
    are invertible, and
  • (iii)
    for all $x,y,z\in X$ we have the exchange laws:
    1. (1)
      $(x\,\underline{\triangleright}\,y)\,\underline{\triangleright}\,(z\,\underline{\triangleright}\,y)=(x\,\underline{\triangleright}\,z)\,\underline{\triangleright}\,(y\,\overline{\triangleright}\,z),$
    2. (2)
      $(x\,\overline{\triangleright}\,y)\,\underline{\triangleright}\,(z\,\overline{\triangleright}\,y)=(x\,\underline{\triangleright}\,z)\,\overline{\triangleright}\,(y\,\underline{\triangleright}\,z),$
    3. (3)
      $(x\,\overline{\triangleright}\,y)\,\overline{\triangleright}\,(z\,\overline{\triangleright}\,y)=(x\,\overline{\triangleright}\,z)\,\overline{\triangleright}\,(y\,\underline{\triangleright}\,z).$
Axiom (ii) is equivalent to the adjacent labels rule, which says that in the ordered quadruple $(x,y,x\,\underline{\triangleright}\,y,y\,\overline{\triangleright}\,x)$, any two neighboring entries (including $(y\,\overline{\triangleright}\,x,x)$ determine the other two. A biquandle is a set $X$ with a choice of biquandle structure.
Example 4.2.
A $\mathbb{Z}[t^{\pm 1},s^{\pm 1}]$-module has a biquandle structure known as an Alexander biquandle defined by
$x\,\underline{\triangleright}\,y=tx+(s-t)y\quad\mathrm{and}\quad x\,\overline{\triangleright}\,y=sx.$
In particular, a choice of units $t,s\in\mathbb{Z}_{n}$ defines an Alexander biquandle structure on $\mathbb{Z}_{n}$.
{lpic}
[…]
[r]125,95;$x\,\overline{\triangleright}\,y=x\,\underline{\triangleright}\,y$lbl[r]125,52;$y\,\overline{\triangleright}\,x=y\,\underline{\triangleright}\,x$lbl[l]138,55;$x$lbl[l]138,95;$y$
[…]
[r]125,35;$x\,\overline{\triangleright}\,y=x\,\underline{\triangleright}\,y$lbl[r]125,-5;$y\,\overline{\triangleright}\,x=y\,\underline{\triangleright}\,x$lbl[l]138,-5;$x$lbl[l]138,35;$y$
Figure 12. Labeling rules at crossings, markers, and singular points.
Let $X$ and $Y$ be biquandles. A function $f:X\rightarrow Y$ is called a biquandle homomorphism if $f({x}\,\underline{\triangleright}\,{y})={f(x)}\,\underline{\triangleright}\,{f(y)}$ and $f({x}\,\overline{\triangleright}\,{y})={f(x)}\,\overline{\triangleright}\,{f(y)}$ for any $x,y\in X.$ We denote the set of all biquandle homomorphisms from $X$ to $Y$ by ${\rm Hom}(X,Y)$. A bijective biquandle homomorphism is called a biquandle isomorphism. Two biquandles $X$ and $Y$ are said to be isomorphic if there is a biquandle isomorphism $f:X\to Y$.
Let $D$ be a singular marked diagram and let $C(D)$, $V(D)$ and $S(D)$ denote the set of all crossings, marked vertices and singular points of $D$, respectively. By a semi-arc of $D$ we mean a connected component of $D\setminus(C(D)\cup V(D)\cup S(D))$.
Definition 4.3.
Let us fix a biquandle $X$, let $D$ be an oriented singular marked diagram of an oriented immersed surface-link, and let $\Lambda$ be the set of components of this diagram. A biquandle coloring ${\mathcal{C}}$ is a mapping ${\mathcal{C}}:\Lambda\to X$ such that around each classical, marked or singular point, the relation shown in Figure 12 holds. These conditions are consistent around each classical, marked or singular point due to the axioms for the biquandle. Denote by ${Col}_{X}(D)$ the set of all colorings of the diagram $D$ with biquandle $X$.
Proposition 4.4.
The biquandle axioms are chosen such that given a biquandle coloring of one side of any $\Gamma$–type move, there is a unique biquandle coloring of the other side of the move with the condition that colors agree on the boundary arcs that leaves the disc where the move is performed.
Proof.
The proof for the moves $\Gamma_{1},\ldots,\Gamma_{3}$ can be found in the existing literature (see [KNS16]), the axioms for the biquandle were motivated to satisfy those Reidemeister moves, for moves $\Gamma_{4},\ldots,\Gamma_{8}$ see for example [JouNel20]. We checked, by standard hand calculations, the proof for the remaining moves, using the diagrams shown in Figures 12, 911.
In particular, checking the validity for the moves $\Gamma_{9}$ and $\Gamma_{9}^{\prime}$ is similar to the case of $\Gamma_{3}$ case. The validity for the moves $\Gamma_{12}a-\Gamma_{12}d$ and $\Gamma_{11}a-\Gamma_{11}d$ are similar to the $\Gamma_{4}$ and $\Gamma_{4}^{\prime}$ cases, because of the fact that $x\,\overline{\triangleright}\,y=x\,\underline{\triangleright}\,y$ and $y\,\overline{\triangleright}\,x=y\,\underline{\triangleright}\,x$ around the singular point in Figure 12.
Ashihara [Ash12] gave a method to calculate the fundamental biquandle $BQ(L)$ of an oriented surface-link from its marked graph diagram. To get a fundamental biquandle $BQ_{S}(L)$ of an oriented singular surface-link we can take the analogous presentation $S\;|\;R\right>$ as the quotient of the free biquandle on the set $S$ (generators associated to the semiarcs of the diagram) by the equivalence relation generated by relations $R$ (see Figure 12) from any diagram for $L$ (see [KKKL18] for more details of the construction). From the previous Proposition the biquandles $BQ_{S}(L_{1})$ and $BQ_{S}(L_{2})$ are isomorphic if $L_{1}$ and $L_{2}$ are equivalent singular surface-links.
Corollary 4.5.
The number of biquandle colorings #$\#Col_{X}(D)$ of an oriented singular marked diagram $D$ is an invariant of an oriented singular surface-link $\mathcal{L}$ presented by $D$, we can denote it therefore by #$\#Col_{X}(\mathcal{L})$ and also called biquandle coloring invariant.
Let us consider the singular surface link $L_{T}$, shown as the singular marked diagram $D_{T}$ in Figure 13. Let a biquandle on $X_{T}=\{1,2,3\}$ be the biquandle with the operation given by the following matrix.
¯$\begin{array}[]{r|rrr}\,\underline{\triangleright}&1&2&3\\ \hline\cr 1&3&1&3\\ 2&2&2&2\\ 3&1&3&1\end{array}\quad\begin{array}[]{r|rrr}\,\overline{\triangleright}&1&2&3\\ \hline\cr 1&3&3&3\\ 2&2&2&2\\ 3&1&1&1\end{array}$
We can enumerate that #$\#Col_{X_{T}}(D_{T})=5$, two of the admissible coloring is shown in Figure 13, the other two admissible colorings are obtained from the previous by exchanging colors $1$ with $3$, and the fifth one is monochromatic, colored by element $2$.
{lpic}
[]michal_032(9cm)
Figure 13. Admissible colorings of the diagram $D_{T}$.